dan Multiplication by t. 8. Find the following Laplace transforms using the formula L[tf(t)] = (-1), (a) [t3e-36] (b) C[(t
Answer #1

dan Now (8) (9) Llt - 3t) = ? Now Lit 814) = 65 d FUN! Now 811) eat, L18!!))= 543 Now Lit3+) = -13° 4 (1) = (- 19453) =H latet)= alltet) - 2 (-1) d (6) = 2.4-16-1863-15 are (s-12 Now , {(t+1) et] = alt et) + Llyet) + 2(dtet) co to +3932 a s-1 2& 1lt larat) = -1-1) (-1) 1 s +4-282 (+4). 1 - 8 (28) 1 I (3²+4)2 1 (2+4) 2 11+(12+) - szy 1844)? So, L[t (3 sinat - 2 carat)21tsint) = 42). da list) = (-2)./ (3471).1 - 8.2(371). 23 (32+1)9 2 as – 45-48. =(-2) (3 + 1 + 25% - 43 _482 (3+1) 4 a - 2 (

If you have any doubt feel free to ask in the comment section.